OpenStudy (anonymous):

which of the following will form the composite function g(f(x)) shown below? G(F(x))=x^3-5 A) F(x)=x^3 and G(x)=x-5 B) F(x)= x and G(x)=x^3 C) F(x)= x-5 and G(x)=x^3 D) F(x)= x^3 and G(x)=-5

OpenStudy (ranga):

Try each choice: A) G(x) = x - 5; F(x) = x^3 G(F(x)) = G(x^3) = x^3 - 5 Try other choices too to make sure there are not more than one correct answer.

OpenStudy (anonymous):

i think its A

OpenStudy (anonymous):

yup it is

OpenStudy (ranga):

alright.

OpenStudy (anonymous):

thanks

OpenStudy (ranga):

yw.